I need help solving this word problem. I have spent the last half an hour trying out different things and cannot get the math right. Please help.
Four hundred people attended a recent concert. Admission was $10 for each adult and $5 for each child. Ticket receipts totaled $3250. How many children and how many adults attended the concert?
All I have is:
People = 400
Adult = $10
Child = $5
Total = $3250.

Answer by rfer(16322)

a+c=400 a=400-c
10a+5c=3250
10(400-c)+5c=3250
4000-10c+5c=3250
-5c=-750
c=150 children
400-150=250 adults
-------------------
250*10=$2500
150*5=$750
Total=$3250